Circular No. 8995

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 T3 = 1892 T1 (BARNARD-BOATTINI) Following a suggestion by M. Meyer (Limburg, Germany) that comet

P/2008 T3 (cf. IAUC 8993) may be identical to the long-lost comet D/1892 T1 = 1892e = 1892 V (Barnard), G. V. Williams (Smithsonian Astrophysical Observatory) has collected 41 astrometric observations from the literature in 1892 and re-reduced them for analysis. The resulting orbital calculations by Williams and by S. Nakano (Sumoto, Japan) confirm the identity. Nakano notes that the comet made 20 revolutions between 1892 and 2008, passing ∼ 0.3–0.4 AU of Jupiter in 1922 Oct., 1934 Aug., and 2005 Aug. The following orbital elements are by Williams, from 90 observations spanning 1892 Oct. 16–2008 Oct. 8 (22 of which are from 1892; mean residual 1′′.4):

Epoch = 1892 Dec. 8.0 TT

T = 1892 Dec. 11.2241 TT  ω = 170.^{o}0322
e = 0.588534  Ω = 208.0355 }2000.0
q = 1.431937 AU  i = 31.2616

a = 3.480083 AU n^{o} = 0.1518167 P = 6.49 years Epoch = 2008 Oct. 21.0 TT

T = 2008 Oct. 25.0966 TT  ω = 181.^{o}4238
e = 0.646433  Ω = 204.1371 }2000.0
q = 1.145243 AU  i = 32.9309

a = 3.239112 AU n^{o} = 0.1690695 P = 5.83 years

G. Sostero, E. Guido, and P. Camilleri report that sixteen 60-s remotely acquired unfiltered CCD images obtained with a 35-cm reflector at Trunkey, N.S.W., on Oct. 8.4 UT show a compact coma ∼ 8′′ in diameter and a tail ∼ 15′′ long in p.a. 255^{o}.

COMET C/2008 R3 (LINEAR) Elliptical orbital elements for this comet (cf. IAUC 8973) from MPEC

2008-T97:

T = 2008 Nov. 22.467 TT  ω = 84.^{o}139
e = 0.89537  Ω = 270.568 }2000.0
q = 1.90883 AU  i = 43.238
a = 18.24389 AU n^{o} = 0.012648 P = 77.92 years
2008 October 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green